Circular No. 6051

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

X-RAY NOVA IN SCORPIUS R. A. Kroeger, J. E. Grove, J. D. Kurfess, W. N. Johnson, and M.

S. Strickman, Naval Research Laboratory, write: “We have further constrained the position determined by BATSE for GRO J1655–40 (IAUC 6046) using the OSSE instrument on the Compton Gamma Ray Observatory. Data from a 1-dimensional scan through the source region constrain the position to be within 0^{o}.11 (90-percent confidence) of the great circle connecting the following points: α = 16^{h}53^{m}, δ = −39^{o}43′ (equinox 2000.0); α = 16^{h}59^{m}, δ = −41^{o}21′. Our limit overlaps with ∼ 30 percent of the original BATSE error circle and is not consistent with the x-ray source 1ES 1649–403 (cf. IAUC 6048). The γ-ray spectrum is well-represented by a power-law model that continues to harden with time, with the following spectral indicies: Aug. 4, −2.7; 9, −2.4. BATSE’s reported spectral index on Aug. 1 was −3.15. Significant source flux is observed to 600 keV.”

NEW VARIABLE STAR IN SAGITTARIUS H. Sato, National Astronomical Observatory of Japan, has reported

the discovery of what was termed a possible nova by Yukio Sakurai (Mito, Japan) on Fuji SG400 film at the following position: α = 17^{h}56^{m}35^{s}.54, δ = −25^{o}13′15′′.7 (equinox 1950.0); giving the following magnitude estimates: May 20 UT, [12.5; June 17, [11.5; July 3, [11:; 28.520, 11.4; Aug. 8.548, 11.4. Further observations suggest that the object is evidently a Mira-type or symbiotic star undergoing an outburst: T. Beers (Michigan State University), W. Brandner and T. Lehmann (European Southern Observatory = ESO), and H. W. Duerbeck (Astronomical Institute Münster) report that a spectrogram (range 360-495 nm), taken on Aug. 10.084 with the ESO 1.52m telescope, shows strong, narrow (FWHM ≤ 120 km/s) emission lines of Hβ to H_{10} (with Hϵ weak or absent), superimposed on a red continuum with traces of TiO bands at 458.4, 462.6, 476.1, and 480.4 nm. Photometry on Aug. 10.161 yields V = 12.41 ± 0.02, B − V = +2.52 ± 0.04, V − R = +1.56 ± 0.02. M. Della Valle (Universitá di Padova) and A. Smette (Groningen) report that spectrograms (range 400-750 nm, resolution 0.8 nm) obtained on Aug. 10.0 at the La Silla 3.6-m telescope (+ EFOSC) reveal an Me spectrum, characterized by TiO absorption bands furrowed by narrow Balmer emission lines; inspection of photographic surveys yields the following magnitudes: 1950 July, B ∼ 19 (Palomar); 1976 May, B = 14–15 (ESO); 1977 June, R ∼ 12 (Palomar).

1994 August 11 Daniel W. E. Green